0 0
Add a comment Improve this question Transcribed image text
Answer #1

ques 13) Let the smaller number be x

Then the large number is 6x.

So, 6x-x=25, or x=5.

Option (3)

So, the value of smaal number is 5 and large number is 30.

ques 14) radius=100 miles. speed = 40 miles/hour.

distance to go=2*100=200 miles.

So, the driver can listen to the signal till 200/40 = 5 hour.

Option (4) is the answer.

ques 15) total distance travelled= 1.5*8+(24/60)*55

=12+22 miles

=34 miles.

Therefore option(2) is the answer

ques 16) option (b) is the answer because we dont know the total earnings he is getting.For that to know, we must find out the no of hours he worked at his job.

ques 17) Total input from Mary=40+8*1=$48

total output = 9*8=$72

Therefore, total profit=$72-$48=$24

so, profit magrin =total profit/total input from Mary=24/48=50%

So, option (D) is the answer.
Answer #2

13)- 5

14)- 5 hrs

15)- 34 miles

16) - the number of hours he works each job

17)- 33 %

If u have any doubt in any part feel free to ask any doubts
Answer #3

13)

a - b = 25

b = a/6

=> a - a/6 = 25

=> 5a/6 = 25

=> a = 30

=>b = 5

14) 100/40 = 2.5 hours

15) Total distance = 8 miles + 55 miles/ 60 minutes * 24 = 30 miles

16) The number of hours he works at each job.

17) ($9 - $6 / $9) * 100 = 33%